the mean lifetime of a sample of 100 light tubes produced by a company is found to be 1,580 hours. Test the hypothesis at 5% level of significance that the mean lifetime of the tubes produced by the company is 1,600 hours with standard deviation of 90 hours


1
Expert's answer
2022-05-24T13:21:00-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=1600"

"H_1:\mu\not=1600"

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a two-tailed test is "z_c = 1.96."

The rejection region for this two-tailed test is "R = \{z:|z|>1.96\}."

The z-statistic is computed as follows:



"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{1580-1600}{90/\sqrt{100}}\approx-2.2222"

Since it is observed that "|z|=2.2222>1.96=z_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value is "p=2P(z<-2.2222)=0.02627," and since "p= 0.02627<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is different than 1600, at the "\alpha = 0.05" significance level.